Determine the ordered pair solution for the given linear equations. state the answer in 2 decimal places.

-x/5-y/7=3 and y/3-x/2=1/11

Answer by rothauserc(4718)

1) -x/5 - y/7 = 3
2) y/3 -x/2 = 1/11
:
Equation 1)'s common denominator is 35, multiply both sides of = by 35
:
3) -7x -5y = 105
:
Equation 2)'s common denominator is 66, multiply both sides of = by 66
:
4) 22y -33x = 6
:
solve equation 3) for x
:
-7x = 5y +105
:
x = -5y/7 -15
:
substitute for x in equation 4)
:
22y -33(-5y/7 -15) = 6
:
22y +165y/7 +495 = 6
:
319y/7 = -489
:
319y = -3423
:
y = -10.73
x = -5(-10.73)/7 -15 = -7.34
:
*****************************************
the ordered pair is (-7.34, -10.73)
*****************************************
:
substitute for x and y in the equation
:
y/3 -x/2 = 1/11
;
-10.73/3 -(-7.34)/2 = 1/11
:
-3.58 +3.67 = 1/11
:
0.09 = 0/09
:
our answer checks :-)
: